OpenStudy (anonymous):

If SuperflyTyGuy works 3/4 of hour on Monday and 6/10 of hour on Tuesday. how much time will he work together?

OpenStudy (superflytyguy):

I worked 3/4 of an on Monday and 6/10 on Tuesday. 6/10 = 12/20 and 3/4 = 15/20 = 27/20 = 1 7/20

OpenStudy (matlee):

81minutes

OpenStudy (anonymous):

Superfly I have lots more questions for you